# Preferential Attachment Trees with Vertex Death: Persistence of the Maximum Degree

### v1

@@ 4.3. Preliminary results
where we recall $I(u)$ from (4.72). In words, $\mathcal{A}^{-}$ consists of all individuals $u$ born up to time $T$ where, for each individual $u$ such that $D^{(u)}=\infty$, its sub-tree is restricted to the first $K$ children and its descendants (if they are born by time $T$). Then, $\mathcal{A}^{+}$ consists of the individuals that are not yet born by time $T$ among the first $I(u)$ children of individuals $u\in\mathcal{A}^{-}$.
With these sets at hand, we define the following events. For $u\in\mathcal{A}^{-}$, set
$\mathcal{C}_{u}:=\{uv:v\in\mathcal{U}_{\infty},v_{1}>K,D^{(uv)}=\infty,\text{ $uv$ catches up to $u$}\},$ (4.78)
and for $u\in\mathcal{A}^{+}$ we also define
$\mathcal{D}^{(u)}:=\{uv:v\in\mathcal{U}_{\infty},D^{(uv)}=\infty,\text{ $uv$ catches up to $V$}\}.$ (4.79)
With these events at hand, we observe that
$\{u\in\mathcal{U}_{\infty}:{}$ $D^{(u)}=\infty,\ \mathcal{B}(u)>\mathcal{B}(T),\text{ $u$ catches up to all individuals in }\mathrm{Pre}(u)\}$ (4.80)
@@
$\subseteq\bigg{(}\bigcup_{\begin{subarray}{c}u\in\mathcal{A}^{-}\\ D^{(u)}=\infty\end{subarray}}\mathcal{C}_{u}\bigg{)}\cup\bigg{(}\bigcup_{u\in\mathcal{A}^{+}}\mathcal{D}^{(u)}\bigg{)}.$
Indeed, any individual $w$ born after time $T$ satisfies that it has an ancestor in $\mathcal{A}^{+}$ or, if not, it has a ancestor $u\in\mathcal{A}^{-}$ such that $D^{(u)}=\infty$ and $w=uv$ with $v_{1}>K$. Then, any such individual $w$ that catches up to all its predecessors and that has an ancestor in $\mathcal{A}^{+}$ needs to catch up to $V$. If it does not have an ancestor in $\mathcal{A}^{+}$ but has an ancestor $u\in\mathcal{A}^{-}$ with $D^{(u)}=\infty$, it needs to catch up to $u$. We now show that each of these sets has a finite size $\mathbb{P}_{\mathcal{S}}$-almost surely.
First, it is clear that $\mathcal{A}^{-}\subseteq\mathrm{BP}(T)$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely, as already mentioned above. Since for each individual in $\mathcal{A}^{-}$ we include only a finite number of children in $A^{+}$, it follows that $\mathcal{A}^{+}$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely as well. It thus suffices to show that $|\mathcal{C}_{u}|$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely for any $u\in\mathcal{A}^{-}$ such that $D^{(u)}=\infty$, and that $|\mathcal{D}^{(u)}|$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely for any $u\in\mathcal{A}^{+}$.
The former claim follows directly from Proposition 4.13 by replacing $D^{(\varnothing)}$ with $D^{(u)}$ and ‘$v$ catches up to $\varnothing$’ by ‘$uv$ catches up to $u$’ in the statement. As a result,
@@
For the latter claim we use that, by its definition, any individual $u\in\mathcal{A}^{+}$ is born strictly after time $\mathcal{B}(VK)=T$. With $\mathcal{F}_{T}$ the $\sigma$-algebra generated by the process up to time $T$, we write for $u\in\mathcal{A}^{+}$,
$\mathbb{E}_{\mathcal{S}}\left[|\mathcal{D}^{(u)}|\,\Big{|}\,\mathcal{F}_{T}\right]=\sum_{v\in\mathcal{U}_{\infty}}\mathbb{P}_{\mathcal{S}}\left(D^{(uv)}=\infty,\ uv\text{ catches up to }V\,\Big{|}\,\mathcal{F}_{T}\right).$ (4.82)
As in the proof of Proposition 4.13, for each $v=v_{1}\cdots v_{m}$ with $m\in\mathbb{N}_{0},v_{1},\ldots,v_{m}\in\mathbb{N}$ we can write the probability as
@@
$\exists j\in\mathbb{N}_{0}:\mathcal{B}(u)+\sum_{\ell=1}^{m}\sum_{i=1}^{v_{\ell}}X(uv_{1}\ldots v_{\ell-1}i)+\sum_{i=1}^{K+j}X(uvi)\leq\mathcal{B}(VK)+\sum_{i=K+1}^{K+j}X(Vi)\,\Bigg{|}\,\mathcal{F}_{T}\Bigg{)}.$
We note that $T=\mathcal{B}(VK)$ and that $\mathcal{B}(u)>T$. The offspring random variables $D^{(uv_{1}\cdots v_{\ell-1})}$ and $D^{(uv)}$ are not contained in the $\sigma$-algebra $\mathcal{F}_{T}$. Also, as an individual $V$ that produces an infinite offspring has been born by time $T$, survival is guaranteed, so that we can omit the conditioning on the event $\mathcal{S}$. By the independence of the offspring and inter-birth random variables, and since $\mathcal{B}(u)>T=\mathcal{B}(VK)$, we obtain the upper bound

### v2

@@ 4.2. Preliminary results
where we recall $I(u)$ from (4.60) and $\mathcal{U}_{K}^{<\infty}$ from (4.61). We think of $\mathcal{A}_{+}$ as the ‘boundary’ of $\mathcal{A}_{-}$ restricted to $\mathcal{U}_{K}^{<\infty}$. That is, $\mathcal{A}_{+}$ consists of all individuals in $\mathcal{U}_{K}^{<\infty}$ that have not been born by time $T$ but whose parent has been born by time $T$. We then have
$\mathcal{U}^{<\infty}\subseteq\mathcal{A}_{-}\cup\bigcup_{\begin{subarray}{c}u\in\mathcal{A}_{-}\\ D^{(u)}=\infty\end{subarray}}\{uv\colon v\in\mathcal{U}_{\infty},v_{1}>K\}\cup\bigcup_{u\in\mathcal{A}_{+}}\{uv\colon v\in\mathcal{U}_{\infty}\}.$ (4.66)
Indeed, let $w\in\mathcal{U}^{<\infty}$. We show that $wj$ is in one of the sets on the right-hand side for each $j\in[D^{(w)}]$ when $D^{(w)}<\infty$ and for each $j\in\mathbb{N}$ when $D^{(w)}=\infty$. Since $\varnothing\in\mathcal{A}_{-}$, this recursively implies the inclusion.
First, suppose that $w\in\mathcal{A}_{-}$. If $D^{(w)}<\infty$, then for any $j\in[D^{(w)}]$ the individual $wj\in\mathcal{A}_{-}$ if $wj$ is born by time $T$, or it is in $\mathcal{A}_{+}$ if it is not born by time $T$ (and so it is included in the set $\{wjv\colon v\in\mathcal{U}_{\infty}\}$). If $D^{(w)}=\infty$, then for any $j\in[K]$ we can repeat the same argument as in the previous case, and for any $j>K$ the individual $wj$ is in $\{wv\colon v\in\mathcal{U}_{\infty},v_{1}>K\}$.
Second, suppose that $w\in\{uv\colon v\in\mathcal{U}_{\infty},v_{1}>K\}$ for some $u\in\mathcal{A}_{-}$ with $D^{(u)}=\infty$. It then directly follows that $wj\in\{uv\colon v\in\mathcal{U}_{\infty},v_{1}>K\}$ for any $j\in\mathbb{N}$ (irrespective of the value of $D^{(w)}$).
Third, suppose that $w\in\{uv\colon v\in\mathcal{U}_{\infty}\}$ for some $u\in\mathcal{A}_{+}$. Again, it directly follows that $wj\in\{uv\colon v\in\mathcal{U}_{\infty}\}$ for any $j\in\mathbb{N}$ (irrespective of the value of $D^{(w)}$).
Using this superset, we then construct a superset for $\mathcal{P}_{2}$. To this end, we observe that $u\in\mathcal{P}_{2}$ implies that $D^{(u)}=\infty$ and that, for any $w_{1}\neq w_{2}\in\text{Pre}(u)$, either $u$ catches up to $w_{1}$ or $u$ catches up to $w_{2}$, as $u$ catches up to all but at most one predecessor. Abbreviating ‘catches up to’ by c.u.t. and using (4.66), we thus have the inclusion
@@
$\mathcal{P}_{2}\subseteq\mathcal{A}_{-}{}$ $\cup\bigcup_{\begin{subarray}{c}u\in\mathcal{A}_{-}\setminus\{V_{2}\}\\ D^{(u)}=\infty\end{subarray}}\{uv\colon v\in\mathcal{U}_{\infty},v_{1}>K,D^{(uv)}=\infty,uv\text{ c.u.t. }u\text{ or }V_{2}\}$ (4.67)
@@
$\cup\bigcup_{u\in\mathcal{A}_{+}}\{uv\colon v\in\mathcal{U}_{\infty},D^{(uv)}=\infty,uv\text{ c.u.t. }V_{1}\text{ or }V_{2}\}.$
Again, this inclusion is clear. By (4.66), we take all individuals in $u\in\mathcal{A}_{-}\setminus\{V_{2}\}$ such that $D^{(u)}=\infty$ and require an individual in the set $\{uv\colon v\in\mathcal{U}_{\infty},v_{1}>K\}$ to catch up to either $u$ or to $V_{2}$. Similarly, we require any individual in the set $\{V_{2}v\colon v\in\mathcal{U}_{\infty},v_{1}>K\}$ to catch up to either $V_{1}$ or $V_{2}$. Finally, for any individual in a set $\{uv\colon v\in\mathcal{U}_{\infty}\}$ for some $u\in\ \mathcal{A}_{+}$, we require they catch up to either $V_{1}$ or $V_{2}$.
It remains to show that each of the sets on the right-hand side of the inclusion are finite $\mathbb{P}_{\mathcal{S}}$-almost surely. First, since $T$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely it follows from Lemma 4.2 that $\mathcal{A}_{-}\subseteq\mathrm{BP}(T)$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely. As $I(u)$ is finite almost surely for each $u\in\mathcal{U}_{\infty}$, it also follows that $\mathcal{A}_{+}$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely. We also have the inclusions
$\{uv{}\colon{}$ $v\in\mathcal{U}_{\infty},v_{1}>K,D^{(uv)}=\infty,uv\text{ c.u.t. }u\text{ or }V_{2}\}$ (4.68)
[…]
$\cup\{uv\colon v\in\mathcal{U}_{\infty},v_{1}>K,D^{(uv)}=\infty,uv\text{ c.u.t. }V_{2}\},$
and
$\{V_{2}v\colon{}$ $v\in\mathcal{U}_{\infty},v_{1}>K,D^{(V_{2}v)}=\infty,V_{2}v\text{ c.u.t. }V_{1}\text{ or }V_{2}\}$ (4.69)
$\subseteq{}$ $\bigcup_{j=1}^{2}\{V_{2}v\colon v\in\mathcal{U}_{\infty},v_{1}>K,D^{(V_{2}v)}=\infty,V_{2}v\text{ c.u.t. }V_{j}\},$
and
$\{uv{}\colon{}$ $v\in\mathcal{U}_{\infty},D^{(uv)}=\infty,uv\text{ c.u.t. }V_{1}\text{ or }V_{2}\}$ (4.70)
$\subseteq{}$ $\bigcup_{j=1}^{2}\{uv\colon v\in\mathcal{U}_{\infty},D^{(uv)}=\infty,uv\text{ c.u.t. }V_{j}\}.$
We thus obtain the desired result if we show that, for any $u\in\mathcal{A}_{-}$ such that $D^{(u)}=\infty$, the sets on the right-hand side of (4.68) are finite $\mathbb{P}_{\mathcal{S}}$-almost surely, the sets on the right-hand side of (4.69) are finite $\mathbb{P}_{\mathcal{S}}$-almost surely, and, for any $u\in\mathcal{A}_{+}$, the sets on the right-hand side of (4.70) are finite $\mathbb{P}_{\mathcal{S}}$-almost surely. We show that this is indeed the case, distinguishing per equation. The proof for each case is similar to the proof of Proposition 4.8, but with slight, case specific, differences.
[…]
$\{uv\colon v\in\mathcal{U}_{\infty},v_{1}>K,D^{(uv)}=\infty,uv\text{ c.u.t. }V_{2}\}\qquad\text{for }u\in\mathcal{A}_{-}\setminus\{V_{2}\}\text{ such that }D^{(u)}=\infty.$ (4.71)
Let $\mathcal{F}_{T}$ be the $\sigma$-algebra generated by $(\mathrm{BP}(t);t\leq T)$. We note that $T$, $\mathcal{A}_{-}$, $\mathcal{A}_{+}$, $V_{1}$, and $V_{2}$ are all measurable with respect to $\mathcal{F}_{T}$. Take $u\in\mathcal{A}_{-}\setminus\{V_{2}\}$ with $D^{(u)}=\infty$. We distinguish between two cases.
Case (i) $u\neq V_{1}$. Similar to the proof of Proposition 4.8, we can write for any $uv$ with $v=v_{1}\cdots v_{m}\in\mathcal{U}_{\infty}$ and with $m\in\mathbb{N}_{0}$, $v_{1}>K$, and $v_{2},\ldots,v_{m}\in\mathbb{N}$,
$\mathbb{P}_{\mathcal{S}}\big(uv{}$ $\text{ catches up to }V_{2}\text{ and }D^{(uv)}=\infty\,\big|\,\mathcal{F}_{T}\big)$ (4.72)
$=\mathbb{P}_{\mathcal{S}}\Bigg({}$ $D^{(uv)}=\infty,D^{(uv_{1}\cdots v_{\ell-1})}\geq v_{\ell}\text{ for all }\ell\in[m],$
$\exists j\in\mathbb{N}_{0}\colon\mathcal{B}(u)+\sum_{\ell=1}^{m}\sum_{i=0}^{v_{\ell}-1}E^{(uv_{1}\cdots v_{\ell-1})}_{i}+\sum_{i=0}^{K+j}E^{(uv)}_{i}\leq\mathcal{B}(V_{2}K)+\!\!\!\sum_{i=K}^{K+j}E^{(V_{2})}_{i}\,\Bigg|\,\mathcal{F}_{T}\Bigg).$
Since $V_{2}$ produces its $K^{\mathrm{th}}$ child at time $T$ by the definition of $T$ and $V_{2}$, it follows that $\mathcal{B}(V_{2}K)=T$. Furthermore, since we suppose that $u\neq V_{1}$, we know that $u$ has produced at most $K$ children (in fact, fewer than $K$ children) by time $T$, so that $\mathcal{B}(uK)\leq T$. We can thus omit $\mathcal{B}(u)+\sum_{i=0}^{K-1}E^{(u)}_{i}=\mathcal{B}(uK)$ and $\mathcal{B}(V_{2}K)$ from the event to obtain the upper bound
@@
$\exists j\in\mathbb{N}_{0}\colon\sum_{i=K}^{v_{1}-1}E^{(u)}_{i}+\sum_{\ell=2}^{m}\sum_{i=0}^{v_{\ell}-1}E^{(uv_{1}\cdots v_{\ell-1})}_{i}+\sum_{i=0}^{K+j}E^{(uv)}_{i}\leq\sum_{i=K}^{K+j}E^{(V_{2})}_{i}\,\Bigg|\,\mathcal{F}_{T}\Bigg).$
We now observe that all random variables in the conditional probability, with the exception of $D^{(u)}$, are independent of $\mathcal{F}_{T}$. Since we know, however, that $D^{(u)}=\infty$ and thus $D^{(u)}\geq v_{1}$ is satisfied, the probability almost surely equals
$\mathbb{P}_{\mathcal{S}}\Bigg({}$ $D^{(uv)}=\infty,D^{(uv_{1}\cdots v_{\ell-1})}\geq v_{\ell}\text{ for all }\ell\in\{2,\ldots,m\},$ (4.74)
$\exists j\in\mathbb{N}_{0}\colon\sum_{i=K}^{v_{1}-1}E^{(u)}_{i}+\sum_{\ell=2}^{m}\sum_{i=0}^{v_{\ell}-1}E^{(uv_{1}\cdots v_{\ell-1})}_{i}+\sum_{i=0}^{K+j}E^{(uv)}_{i}\leq\sum_{i=K}^{K+j}E^{(V_{2})}_{i}\Bigg)$.
By bounding $\mathbb{P}_{\mathcal{S}}\left(\mathcal{E}\right)\leq\mathbb{P}\!\left(\mathcal{E}\right)/\mathbb{P}\!\left(\mathcal{S}\right)$ for any event $\mathcal{E}$, using the independence of the offspring and inter-birth random variables, and applying Lemma 4.3 (where we note that the condition in (4.8) is satisfied for any $\lambda>0$ by choosing $K$ large enough, see Remark 4.6), we can bound the probability from above by
$\frac{\mathbb{P}\!\left(D=\infty\right)}{\mathbb{P}\!\left(\mathcal{S}\right)}{}$ $\prod_{i=K}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{K-1}E_{i}\bigg)\right]\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=K}^{v_{1}-1}E_{i}\bigg)\right]$ (4.75)
$\times\prod_{\ell=2}^{m}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{v_{\ell}-1}E_{i}\bigg)\right].$
Since the inter-birth random variables in the final two terms on the first line are independent, we can combine them into a single expected value. By also bounding $\mathbb{P}\!\left(D=\infty\right)\leq\mathbb{P}\!\left(D\geq v_{1}\right)$, we thus obtain the upper bound
$\frac{1}{\mathbb{P}\!\left(\mathcal{S}\right)}\prod_{i=K}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]\prod_{\ell=1}^{m}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{v_{\ell}-1}E_{i}\bigg)\right].$ (4.76)
We set $c=\mathbb{P}\!\left(\mathcal{S}\right)^{-1}$. By summing over all $uv$ with $v\in\mathcal{U}_{\infty}$ such that $v_{1}>K$, we thus obtain
$\mathbb{E}_{\mathcal{S}}\Big[{}$ $|\{uv\colon v\in\mathcal{U}_{\infty},v_{1}>K,uv\text{ catches up to }V_{2}\text{ and }D^{(uv)}=\infty\}|\,\Big|\,\mathcal{F}_{T}\Big]$ (4.77)
$\leq{}$ $\sum_{m=1}^{\infty}\sum_{v_{1}=K+1}^{\infty}\sum_{v_{2}=1}^{\infty}\cdots\sum_{v_{m}=1}^{\infty}c\prod_{i=K}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]\prod_{\ell=1}^{m}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=1}^{v_{\ell}}E_{i}\bigg)\right]$
$\leq c\prod_{i=K}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]\sum_{m=1}^{\infty}\Bigg(\sum_{v=1}^{\infty}\mathbb{P}\!\left(D\geq v\right)\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{v-1}E_{i}\bigg)\right]\Bigg)^{m},$
where in the last step we interchanged summation and products and extended the range of the sum over $v_{1}$ to all $v_{1}\in\mathbb{N}$. We observe that the inner sum on the last line equals $\widehat{\mu}(\lambda)$, so that by Assumption (F-L) we can take $\lambda>0$ large enough such that $\widehat{\mu}(\lambda)<1$, as follows from (4.2). The upper bound is thus finite, from which it follows that
$\mathbb{P}_{\mathcal{S}}\left(\{uv\colon v\in\mathcal{U}_{\infty},v_{1}>K,uv\text{ catches up to }V_{2}\text{ and }D^{(uv)}=\infty\}|<1\,\Big|\,\mathcal{F}_{T}\right)=1,\quad\mathbb{P}_{\mathcal{S}}\text{-a.s.}$ (4.78)
Case (ii) $u=V_{1}$. In this case, let
$\mathcal{T}_{T}(V_{1})\coloneq\{V_{1}v\colon v\in\mathcal{U}_{\infty},v_{1}>K,\mathcal{B}(V_{1}v)\leq T\}$ (4.79)
be the sub-tree rooted at $V_{1}$, restricted to the children $V_{1}(K+1),V_{1}(K+2),\ldots$ and their descendants that are born by time $T$. It is clear that $\mathcal{T}_{T}(V_{1})\subset\mathrm{BP}(T)$, so that it is finite $\mathbb{P}_{\mathcal{S}}$-almost surely, and that it is $\mathcal{F}_{T}$ measurable. For $w\in\mathcal{T}_{T}(V_{1})$, we define
$K(w)\coloneq\max\{0\leq k\leq D^{(w)}:wk\in\mathcal{T}_{T}(V_{1})\}$ (4.80)
to be the ‘last’ child of $w$ that is included in $\mathcal{T}_{T}(V_{1})$. We then have the inclusion
$\{V_{1}v:{}$ $v\in\mathcal{U}_{\infty},v_{1}>K,D^{(V_{1}v)}=\infty,V_{1}v\text{ c.u.t.\ }V_{2}\}$ (4.81)
$\subseteq\mathcal{T}_{T}(V_{1})\cup\bigcup_{w\in\mathcal{T}_{T}(V_{1})}\{wv:v\in\mathcal{U}_{\infty},v_{1}>K(w),D^{(wv)}=\infty,wv\text{ c.u.t.\ }V_{2}\}.$
As a result, we have
$\mathbb{E}_{\mathcal{S}}{}$ $\Big[|\{V_{1}v:v\in\mathcal{U}_{\infty},v_{1}>K,\mathcal{B}(V_{1}v)>T,D^{(V_{1}v)}=\infty,V_{1}v\text{ c.u.t.\ }V_{2}\}|\,\Big|\,\mathcal{F}_{T}\Big]$ (4.82)
$\leq{}$ $|\mathcal{T}_{T}(V_{1})|+\!\!\!\!\sum_{w\in\mathcal{T}_{T}(V_{1})}\!\!\!\!\!\!\mathbb{E}_{\mathcal{S}}\left[|\{wv:v\in\mathcal{U}_{\infty},v_{1}>K(w),D^{(wv)}=\infty,wv\text{ c.u.t.\ }V_{2}\}|\,\Big|\,\mathcal{F}_{T}\right].$
As already concluded, the first term on the right-hand side is finite $\mathbb{P}_{\mathcal{S}}$-almost surely. Each expected value in the sum on the right-hand side can be written as
$\sum_{m=1}^{\infty}\sum_{v_{1}=K(w)+1}^{D^{(w)}}\sum_{v_{2}=1}^{\infty}\cdots\sum_{v_{m}=1}^{\infty}\mathbb{P}_{\mathcal{S}}\left(D^{(wv)}=\infty,wv\text{ c.u.t.\ }V_{2}\,\Big|\,\mathcal{F}_{T}\right),$ (4.83)
and each such probability equals
$\mathbb{P}_{\mathcal{S}}\bigg({}$ $\exists j\in\mathbb{N}_{0}\colon\mathcal{B}(wK(w))+\!\!\!\!\!\sum_{i=K(w)}^{v_{1}-1}\!\!\!\!E^{(w)}_{i}+\sum_{\ell=2}^{m}\sum_{i=0}^{v_{\ell}-1}E^{(wv_{1}\cdots v_{\ell-1})}_{i}+\!\sum_{i=0}^{K+j}\!E^{(wv)}_{i}\!\leq\mathcal{B}(V_{2}K)+\!\sum_{i=K}^{K+j}\!E^{(V_{2})}_{i}$, (4.84)
$D^{(wv)}=\infty,D^{(wv_{1}\cdots v_{\ell-1})}\geq v_{\ell}\text{ for all }\ell\in[m]\,\bigg|\,\mathcal{F}_{T}\bigg).$
We observe that $\mathcal{B}(w(K(w)+1))=\mathcal{B}(wK(w))+E^{(w)}_{K(w)}>T=\mathcal{B}(V_{2}K)$ by the definition of $K(w)$, so that we obtain the upper bound
$\mathbb{P}_{\mathcal{S}}\bigg({}$ $\exists j\in\mathbb{N}_{0}:\!\!\!\!\sum_{i=K(w)+1}^{v_{1}-1}\!\!\!\!E^{(w)}_{i}+\sum_{\ell=2}^{m}\sum_{i=0}^{v_{\ell}-1}E^{(wv_{1}\cdots v_{\ell-1})}_{i}+\!\sum_{i=0}^{K+j}\!E^{(wv)}_{i}\!\leq\sum_{i=K}^{K+j}\!E^{(V_{2})}_{i}$, (4.85)
$D^{(wv)}=\infty,D^{(wv_{1}\cdots v_{\ell-1})}\geq v_{\ell}\text{ for all }\ell\in[m]\,\bigg|\,\mathcal{F}_{T}\bigg).\textbf{}$
Only the random variables $D^{(w)}$, $K(w)$, and $V_{2}$ depend on $\mathcal{F}_{T}$. Also, the random variable $K(w)$ is independent of the inter-birth times $E^{(w)}_{i}$ for $i\geq K(w)+1$. As a result, we can apply Lemma 4.3 (where we note that the condition in (4.8) is satisfied for any $\lambda>0$ by choosing $K$ large enough, see Remark 4.6) to obtain the upper bound
$\frac{\mathbb{P}\!\left(D=\infty\right)}{\mathbb{P}\!\left(\mathcal{S}\right)}{}$ $\mathbbm{1}_{\{D^{(w)}\geq v_{1}\}}\prod_{i=K}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{K-1}E_{i}\bigg)\right]\prod_{i=K(w)+1}^{v_{1}-1}\mathbb{E}\left[\mathrm{e}^{-\lambda E_{i}}\right]$ (4.86)
$\times\prod_{\ell=2}^{m}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{v_{\ell}-1}E_{i}\bigg)\right].$
Substituting this upper bound in (4.83), we thus arrive at the upper bound
$\mathbb{E}_{\mathcal{S}}\Big[{}$ $|\{wv:v\in\mathcal{U}_{\infty},v_{1}>K(w),D^{(wv)}=\infty,wv\text{ c.u.t.\ }V_{2}\}|\,\Big|\,\mathcal{F}_{T}\Big]$ (4.87)
$\leq{}$ $\sum_{m=1}^{\infty}\sum_{v_{1}=K(w)+1}^{D^{(w)}}\sum_{v_{2}=1}^{\infty}\cdots\sum_{v_{m}=1}^{\infty}\frac{\mathbb{P}\!\left(D=\infty\right)}{\mathbb{P}\!\left(\mathcal{S}\right)}\prod_{i=K}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{K-1}E_{i}\bigg)\right]$
$\times\prod_{i=K(w)+1}^{v_{1}-1}\mathbb{E}\left[\mathrm{e}^{-\lambda E_{i}}\right]\prod_{\ell=2}^{m}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{v_{\ell}-1}E_{i}\bigg)\right].$
Switching summation and products yields
$\frac{\mathbb{P}\!\left(D=\infty\right)}{\mathbb{P}\!\left(\mathcal{S}\right)}\prod_{i=K}^{\infty}{}$ $\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{K-1}E_{i}\bigg)\right]$ (4.88)
$\times\sum_{v_{1}=K(w)+1}^{D^{(w)}}{}$ $\prod_{i=K(w)+1}^{v_{1}-1}\!\!\!\mathbb{E}\left[\mathrm{e}^{-\lambda E_{i}}\right]\sum_{m=1}^{\infty}\Bigg(\sum_{v=1}^{\infty}\mathbb{P}\!\left(D\geq v\right)\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{v-1}E_{i}\bigg)\right]\Bigg)^{m-1}.$
As in Case (i), we can write the inner sum as $\widehat{\mu}(\lambda)$ and by Assumption (F-L) we can choose $\lambda>0$ such that $\widehat{\mu}(\lambda)<1$ (see the proof of Lemma 4.2). We then observe that
$\sum_{v_{1}=K(w)+1}^{D^{(w)}}\prod_{i=K(w)+1}^{v_{1}-1}\!\!\!\!\!\!\mathbb{E}\left[\mathrm{e}^{-\lambda E_{i}}\right]$ $\leq\sum_{v_{1}=K(w)+1}^{\infty}\prod_{i=K(w)+1}^{v_{1}-1}\!\!\!\!\!\mathbb{E}\left[\mathrm{e}^{-\lambda E_{i}}\right]$ (4.89)
$\leq\mathbb{P}\!\left(D=\infty\right)^{-1}\!\!\!\prod_{i=0}^{K(w)}\!\mathbb{E}\left[\mathrm{e}^{-\lambda E_{i}}\right]^{-1}\sum_{v=1}^{\infty}\mathbb{P}\!\left(D\geq v\right)\prod_{i=1}^{v-1}\mathbb{E}\left[\mathrm{e}^{-\lambda E_{i}}\right]$
$=\mathbb{P}\!\left(D=\infty\right)^{-1}\widehat{\mu}(\lambda)\prod_{i=1}^{K(w)}\mathbb{E}\left[\mathrm{e}^{-\lambda E_{i}}\right]^{-1}.$
This upper bound is $\mathbb{P}_{\mathcal{S}}$-almost surely finite for any $w\in\mathcal{T}_{T}$, since $K(w)\leq K$ by the definition of $T$ and $V_{1},V_{2}$ and since we can take $\lambda>0$ such that $\widehat{\mu}(\lambda)<\infty$ by Assumption (F-L). As a result, since $|\mathcal{T}_{T}(V_{1})|$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely, the upper bound in (4.82) is finite $\mathbb{P}_{\mathcal{S}}$-almost surely. We thus conclude that sets on the left-hand side of (4.68) are finite $\mathbb{P}_{\mathcal{S}}$-almost surely conditionally on $\mathcal{F}_{T}$, for any $u\in\mathcal{A}_{-}\setminus\{V_{2}\}$ .